The tv costs $320. This is $40 more than four times what the old is worth let "w" represent what its worth. and write an equation
1
Expert's answer
2012-12-12T09:15:13-0500
let w be what the old is worth
then 4w is four time. As it is said, if you add 40 to 4w youget 320$, price of new tv
so 4w+40=320
hence
w=(320-40)/4 = 70$